### v3

# Scalable Lattice Sampling using Factorized Generative Models

Abstract
Boltzmann distributions over lattices are pervasive in Computational Physics. Sampling them becomes increasingly difficult with increasing lattice-size, especially near critical regions, e.g., phase transitions in statistical systems and continuum limits in lattice field theory. Machine learning-based methods, such as normalizing flows, have demonstrated the ability to alleviate these issues for small lattices. However, scaling to large lattices is a major challenge. We present a novel approach called Parallelizable Block Metropolis-within-Gibbs (PBMG) for generating samples for any lattice model. It factorizes the joint distribution of the lattice into local parametric kernels, thereby allowing efficient sampling of very large lattices. We validate our approach on the XY model and the Scalar $\phi^{4}$ theory. PBMG achieves high acceptance rates and less correlated samples, and the observable statistics estimated from the samples match the ground truth. Moreover, PBMG significantly speeds up inference for large lattices as compared to HMC and plain MCMC algorithms.
keywords:
Markov Chain Monte Carlo, Metropolis-Hastings, Lattice models, Normalizing flows, Generative machine learning
Lattice models in Physics are mathematical models that contain information in the form of a lattice $\bm{\phi}$ characterized by Boltzmann distribution defined with the help of a Hamiltonian $H(\bm{\phi})$ or action $s(\bm{\phi})$.
$p(\bm{\phi};\bm{\theta})\propto e^{-H(\bm{\phi};\bm{\theta})}$ (1)
Each lattice site is a random variable that could be binary as in the Ising model, circular as in the XY model or real-valued as in scalar $\phi^{4}$ theory. The statistical properties of these lattices vary with parameters $\bm{\theta}$, showing drastic changes near certain regions of $\bm{\theta}$ called critical regions. Although these distributions are known only up to a normalizing constant, they can be sampled using statistical methods such as MCMC (Markov Chain Monte Carlo) [1, 2]. These methods provide convergence guarantees that the samples represent the target distribution $p(\bm{\phi};\bm{\theta})$. But these methods could be quite inefficient, especially in the critical regions. In the critical regions, the correlation between successive samples of the Markov chain (characterized by the autocorrelation time of the chain) becomes very large. This phenomenon is called critical slowing down [3, 4]. Certain algorithms, such as, Swendsen-Wang [5], Wolff [6], worm [7], loop [8], directed loop [9] and HMC (Hamiltonian Monte Carlo) [10], make global MCMC updates to tackle the problem of critical slowing down. However, these methods have not been addressed for large lattices.
Generative machine learning (ML) methods are emerging as promising tools to sample Boltzmann distributions [11]. ML methods, such as Gaussian mixture models (GMM) and normalizing flows (NF), providing exact model probabilities, can be used in Metropolis-Hastings (MH) algorithm to propose samples which are accepted or rejected. Such algorithms provide convergence guarantees too. NF-based methods have been used for $\phi^{4}$ theory[12, 13, 14, 15, 16, 17, 18, 19, 20] and Gauge theory[21, 22, 23, 24, 25, 26, 27, 28, 29] in lattice field theory, and for general Monte Carlo sampling [30]. Self-learning Monte Carlo [31], [32] and Restricted Boltzmann Machines [33] also aim at efficient sampling of lattices using ML methods.
On the other hand, there are ML approaches that do not provide exact model probability, and hence, can not be used in the MH algorithm. These approaches are prone to biases. These include likelihood-free methods, such as, generative adversarial networks [34, 35, 36, 26] and likelihood-based methods, such as, variational autoencoders (VAEs) [37] and Boltzmann generators [38]. These methods produce one-shot samples, as opposed to Markov chains.
To address critical slowing down, some ML methods [39] train the generative model for every $\bm{\theta}$. Since training samples are not available for every $\bm{\theta}$, Reverse Kullback Leibler divergence (RKLD) based learning is used to train the model for a desired $\bm{\theta}$. This approach could be unstable. As an alternative, the use of conditional generative models is a promising approach [34, 40, 41], where the model learns from samples in non-critical regions and extrapolates or interpolates to critical regions.
In other words, do these approaches work efficiently and efficaciously with the increase in lattice size and for different parameters $\bm{\theta}$ (such as quark masses) [42]? Larger lattices provide a finer spatial resolution, allowing one to examine short-distance behaviours more accurately. They also help mitigate finite-size effects, ensuring that simulation results remain reliable and closer to the infinite-volume limit. It facilitates the study of phenomena like confinement, chiral symmetry breaking, and the QCD phase transition, which require the exploration of large-scale structures.
The major hurdle for scalability of learning-based approaches seems to be the joint modeling of the entire lattice. We propose Parallelizable Block Metropolis-within-Gibbs (PBMG) algorithm that factorizes the joint distribution of the lattice and samples in a Metropolis-within-Gibbs setting, with proposals generated by an ML model conditioned on $\bm{\theta}$. Gibbs algorithm has previously been used with ML models to sample large graphs [43], albeit not for sampling from explicit distribution functions.
We present the PBMG method for sampling from conditional Boltzmann distributions over large lattices. We use conditional generative models, e.g., conditional GMMs and conditional NFs, to efficiently sample in critical regions. The conditional generative models are generally trained from given samples. However, to avoid generating ground truth samples for large lattices, we use RKLD-based training. The problem of mode collapse is avoided in PBMG because the local distributions are much simpler (uni-modal or very few modes) as compared to the joint distributions.
To validate the proposed approach, we apply it to 2-D lattices, namely, the XY model from statistical Physics and the scalar $\phi^{4}$ model from lattice field theory. In the XY model, each lattice site is a spin value (angle) that is locally dependent on the immediate neighbors. The lattice distribution is conditioned on temperature $T$ and exhibits phase transition with respect to $T$, where the magnetic susceptibility diverges. We apply PBMG with conditional rational quadratic spline (RQS) flows for XY lattices [44]. In the scalar $\phi^{4}$ theory, each lattice site is a real number locally dependent on the immediate neighbors. The lattice distribution is conditioned on parameters $\lambda$ and $m^{2}$. We study phase transition with respect to $\lambda$, where susceptibility diverges. We apply PBMG with conditional GMMs, conditioned on $\lambda$ and $m^{2}$, for $\phi^{4}$ theory.
The contributions of this work are multifaceted:
  • The introduction of the PBMG algorithm, offering a robust and efficient means to sample Boltzmann distributions over lattices.
  • The pioneering factorization of the joint distribution into local components, laying the foundation for scalable sampling in high-dimensional lattice systems.
  • Proposed an inference algorithm that samples from all the local distributions in a parallel fashion to propose the final lattice sample thereby cutting down significantly on the computational cost.
  • The strategic utilization of RKLD optimization to alleviate the need for ground truth samples.
1 Parallelizable Block Metropolis-within-Gibbs (PBMG)
Consider an $N$-dimensional probability distribution $p(\phi_{1},\phi_{2},\dots,\phi_{N})$. For a lattice, $N$ is the number of lattice sites and $\phi_{i}$ is the random variable (field, dipole, etc.) at site $i$. We partition these sites into $G$ partitions such that the distribution of a site $i$ in a partition $g$, conditioned on all sites $j\notin g$, is independent of all sites $i^{\prime}\in g\setminus i$.
For implementing MCMC in general state-spaces, one requires to construct a Markov chain transition kernel $p(\phi_{i}|\phi_{j\neq i})$ that keeps the target distribution $p(\phi_{1},\phi_{2},\dots,\phi_{N})$ invariant, and is ergodic for this distribution. Such kernels can also be combined via composition. Keeping this in mind, let $K_{i}$ be a transition kernel that updates the site $i\in g$, keeping all other sites of the lattice the same. Then the combined kernel that changes all the sites in the partition $g$ is
$K_{g}=\prod_{i\in g}K_{i}$ (2)
and the overall kernel for updating all the sites in a lattice is
$K=\prod_{g}K_{g}\,.$ (3)
Here, $\prod$ is the iterative composition function. Such a composition is commonplace in deterministic scan Gibbs samplers. The advantage of partitioning is that all the sites in the same partition can be sampled simultaneously, thereby making the process faster. The term “Parallelizable Block” in our algorithm’s nomenclature originates from this inherent capacity to facilitate parallel sampling. Moreover, each kernel $K_{i}$ need not be conditioned on all the sites outside the partition $g$, but only a small number of sites in a local neighbourhood of the site $i$.
Every site-kernel $K_{i}$ for each $i\in g$ and for every partition $g$ is a Metropolis-within-Gibbs kernel, which means that each site-kernel $K_{i}$ is a Gibbs kernel with a Metropolis-Hastings accept-reject step. Metropolis-within-Gibbs algorithms, also known as conditional Metropolis-Hastings, are well studied in the MCMC literature [45].
Let $\bm{\phi}_{-g}$ denote the set of random variables at all the lattice sites excluding the ones in the partition $g$ and $\bm{\psi}$ denote given lattice parameters (e.g., temperature, coupling parameters). Let us define $q(\phi^{(t)}_{i}|\bm{\phi}_{-g},\bm{\psi};\bm{\theta})$ as the parametric proposal distribution, parameterized by $\bm{\theta}$ and conditioned on $\bm{\phi}_{-g}$ and $\bm{\psi}$. Then, the acceptance probability $\alpha_{K_{i}}$ for any site-kernel $K_{i}$ $\forall$ $i\in g$ is
$\alpha_{K_{i}}$ $=\frac{p(\phi^{(t+1)}_{i}|\bm{\phi}_{-g},\bm{\psi})}{p(\phi^{(t)}_{i}|\bm{\phi}_{-g},\bm{\psi})}\cdot\frac{q(\phi^{(t)}_{i}|\phi^{(t+1)}_{i},\bm{\phi}_{-g},\bm{\psi};\bm{\theta})}{q(\phi^{(t+1)}_{i}|\phi^{(t)}_{i},\bm{\phi}_{-g},\bm{\psi};\bm{\theta})}$ (4)
We design a proposal such that,
$q(\phi^{(t+1)}_{i}|\phi^{(t)}_{i},\bm{\phi}_{-g},\bm{\psi};\bm{\theta})=q(\phi^{(t+1)}_{i}|\bm{\phi}_{-g},\bm{\psi};\bm{\theta})$ (5)
i.e., a proposal such that the sampling of the corresponding component is independent of the previous value of the component. Such a proposal is called an independent proposal w.r.t. a component or a conditionally independent proposal. Furthermore, we aim to obtain a high acceptance rate from the proposal. A conditionally independent proposal with a high acceptance rate would ensure that consecutive samples are very much different from each other which would indirectly lower the autocorrelation within the chain.
The maximum acceptance rate possible i.e., an acceptance rate equal to 1 will be achieved when $q(\phi^{(t)}_{i}|\bm{\phi}_{-g},\bm{\psi};\bm{\theta})$ is exactly the same as $p(\phi^{(t)}_{i}|\bm{\phi}_{-g},\bm{\psi})$. This reduces our goal to design(or learn) a proposal that could sample from the true conditional distribution as closely as possible. In order to achieve this goal, we use methods like Normalizing Flows and Gaussian Mixture Models in generative machine learning. To decide which method to use, we need to first analyze the true conditional distribution for a single component in the two PDFs. In the next two sections, we apply the PBMG method to the XY model and the Scalar $\phi^{4}$ theory. In each section, we first analyze the true conditional distribution for a single component in the corresponding PDF. Subsequently, we discuss the structure of the corresponding proposal distribution, the details of the training procedure and the inference procedure.
2 Application to the XY Model
2.1 Target distribution
The Hamiltonian for the XY model is
$H(\bm{\phi})=$ $-\frac{1}{2}\sum_{i,j}\big{[}\cos(\phi_{i,j}-\phi_{i+1,j})+\cos(\phi_{i,j}-\phi_{i,j+1})$
$+\cos(\phi_{i,j}-\phi_{i-1,j})+\cos(\phi_{i,j}-\phi_{i,j-1})\big{]}$ (6)
Here, $\phi_{i,j}$ is the angular random variable with range $[0,2\pi)$ at the lattice site with coordinates $(i,j)$. Here, we have used periodic boundary conditions for the lattice. The lattice distribution at a given temperature $T\in\mathbb{R}$ is
$p(\bm{\phi};T)\propto e^{-\frac{H(\bm{\phi})}{T}}$ (7)
The local Hamiltonian for the $(i,j)$th component $\phi_{i,j}$ of the lattice vector $\bm{\phi}$ is
$H(\phi_{i,j})=$ $-\big{[}\cos(\phi_{i,j}-\phi_{i+1,j})+\cos(\phi_{i,j}-\phi_{i,j+1})$
$+\cos(\phi_{i,j}-\phi_{i-1,j})+\cos(\phi_{i,j}-\phi_{i,j-1})\big{]}$ (8)
We see that the Hamiltonian of the $(i,j)$th component depends only on the components of the four nearest neighbours denoted by $n(i,j)=\{(i+1,j),(i,j+1),(i-1,j),(i,j-1)\}$. Therefore the conditional distribution of $\phi_{i,j}$ given the four nearest neighbour components and temperature is
$p\left(\phi_{i,j}|\{\phi_{l,m}:(l,m)\in n(i,j)\},T\right)=p(\phi_{i,j}|\mathbf{v}_{i,j})\propto e^{-\frac{H(\phi_{i,j})}{T}}$ (9)
The above conditional distribution is our target distribution. Here, $\mathbf{v}_{i,j}=(\phi_{i+1,j},\phi_{i,j+1},\phi_{i-1,j},\phi_{i,j-1},T)$ is the $5$x1111 condition vector corresponding to the site $(i,j)$ which consists of the four nearest neighbour components and the temperature. For this model, we have divided the lattice into two partitions $g_{0}$ and $g_{1}$.
$g_{k}=\{(i,j):(i+j)\%2=k\};k=0,1$ (10)
Figure 1: Partitioning used for a lattice in the XY model. Here, the colors white and black represent the two partitions.
2.2 Modeling the Proposal distribution
Figure 2: Proposal distribution for the XY model
We use Normalizing Flows to model the proposal distribution $q(\phi_{i,j}|\mathbf{v}_{i,j};\bm{\theta})$. $p_{Z}(z|\mathbf{v}_{i,j};\bm{\theta}_{B})$ is the base distribution and $f(z;\bm{\theta}_{R})$ is the invertible transformation used in the Normalizing Flow. Here, $\bm{\theta}=\{\bm{\theta_{B}},\bm{\theta_{R}}\}$. The parameters of $NN_{1}$ is denoted as $\bm{\theta}_{B}$ and the combined set of parameters of $NN_{2},NN_{3}$ and $NN_{4}$ is denoted as $\bm{\theta}_{R}$. The condition vector that is input to all the neural networks is $\mathbf{v}_{i,j}$.
Using the change of variables formula,
$q(\phi_{i,j}|\mathbf{v}_{i,j};\bm{\theta})=p_{Z}\left(f^{-1}(\phi_{i,j};\bm{\theta}_{R})|\mathbf{v}_{i,j};\bm{\theta}_{B}\right)\left|\det{\left(\frac{\partial f^{-1}(\phi_{i,j};\bm{\theta}_{R})}{\partial\phi_{i,j}}\right)}\right|$ (11)
We use Rational Quadratic Splines (RQS) as the transform $f$. An RQS flow is used to map a fixed interval onto another fixed interval. The intervals are partitioned into $K$ bins where each bin is characterized by a rational-quadratic function that increases monotonically. Furthermore, these functions are characterized by $K+1$ coordinates (referred to as knots) representing boundary values, $K-1$ derivative values $\mathbf{s}_{knots}$ at internal knots and the widths $\mathbf{w}_{bins}$ and the heights $\mathbf{h}_{bins}$ of the $K$ bins. For details on RQS flow, one can refer to [44]. The coefficients governing the RQS transformation’s behaviour are adaptable and learned through neural networks. The base distribution is chosen to be uniform, and the interval of the uniform distribution $w_{int}$ is also learnable. We use four different neural networks to learn these parameters, as shown in figure 2.
The parameters of the RQS flow are obtained from the three neural networks $NN_{2},NN_{3},NN_{4}$, and the interval width of the base distribution is obtained from $NN_{1}$. The architecture of each of the four neural networks is the same except for the output dimension, which is dependent on the output parameter. Each neural network has one hidden layer with 1000 neurons and a ReLU activation function. The activations used in the output layers are different. The outputs of $NN_{1}$ and $NN_{4}$ are passed through a Softplus activation function to give the interval width of the base distribution and the $K-1$ slopes at the intermediate knots, respectively. The outputs of $NN_{2}$ and $NN_{3}$ are passed through a $w_{int}$-scaled Softmax activation and a $2\pi$-scaled Softmax activation function to give the widths and the heights of the $K$ bins respectively. We have taken $K$ to be 8.
2.3 Training and Inference Procedure
The loss function used in the training procedure is the expected value of the KL divergence between the proposal $q(\phi_{i,j}|\mathbf{v}_{i,j};\bm{\theta})$ and the target $p(\phi_{i,j}|\mathbf{v}_{i,j})$ i.e., the true conditional distribution over all possible values of the condition vector $\mathbf{v}_{i,j}$ during training. The first four components of $\mathbf{v}_{i,j}$ lie in the interval $[0,2\pi]$ and the last component $T\in[0.13,2.05]$. We will, therefore, sample $\mathbf{v}_{i,j}$ from $p_{v}(\mathbf{v}_{i,j})=$ Unif$([0,2\pi]^{4}\times[0.13,2.05])$ to calculate the expectation.
$\begin{split}\mathcal{L}&=\mathbb{E}_{\mathbf{v}_{i,j}\sim p_{v}(\mathbf{v}_{i,j})}\left[\mathbb{E}_{z\sim p_{Z}(z|\mathbf{v}_{i,j};\bm{\theta}_{B})}\left[\log{q(\phi_{i,j}|\mathbf{v}_{i,j};\bm{\theta})-\log{p(\phi_{i,j}|\mathbf{v}_{i,j})}}\right]\right]\\ &=\mathbb{E}_{\mathbf{v}_{i,j}\sim p_{v}(\mathbf{v}_{i,j})}[\mathbb{E}_{z\sim p_{Z}(z|\mathbf{v}_{i,j};\bm{\theta}_{B})}[\log{p_{Z}(z|\mathbf{v}_{i,j};\bm{\theta}_{B})}+\log{|\det{J_{f}(z|\mathbf{v}_{i,j};\bm{\theta}_{R})}|^{-1}}\\ &-\log{p\left(f(z;\bm{\theta}_{R})|\mathbf{v}_{i,j}\right)}]]\end{split}$ (12)
The Monte Carlo approximation can be used to estimate the above expectation as follows
$\mathcal{L}\approx\frac{1}{n}\cdot\frac{1}{N}\sum_{r=1}^{n}\sum_{k=1}^{N}[\log{p_{Z}(z_{k}|(\mathbf{v}_{i,j})_{r};\bm{\theta}_{B})}+\log{|\det{J_{f}(z_{k}|(\mathbf{v}_{i,j})_{r};\bm{\theta}_{R})}|^{-1}}\\ -\log{p\left(f(z_{k};\bm{\theta}_{R})|(\mathbf{v}_{i,j})_{r}\right)}]$ (13)
We sample from $p_{v}(\mathbf{v}_{i,j})$ to generate 10,000 samples of $\mathbf{v}_{i,j}$ and use this as the training set. We use the Adam optimizer with default hyperparameters and a cosine decay schedule for the learning rate with an initial learning rate of 0.0005 and 20,000 decay steps. We perform validation by calculating the average acceptance rate for 24 random temperatures in the range [0.13, 2.05] and stop training when we achieve an average acceptance rate of around 85%. Our model gets trained quickly, taking less than an hour to train on a low-end single GPU machine.
The procedure for MCMC sampling using PBMG-XY is briefed in the algorithm given below. Here, $\mathbf{V}_{g}=[\mathbf{v}_{i,j}]_{(i,j)\in g}$.
Initialize: $\bm{\phi}$ as $\bm{\phi}^{(0)}$
for $t=0$ to $Num-1$ do
       Calculate $\mathbf{V}_{g_{0}}$ at timestep $t$
       Condition $K_{g_{0}}$ on $\mathbf{V}_{g_{0}}$ and obtain from it $\phi^{(t+1)}_{i,j}\forall(i,j)\in g_{0}$
       Calculate $\mathbf{V}_{g_{1}}$at timestep $t$
       Condition $K_{g_{1}}$ on $\mathbf{V}_{g_{1}}$ and obtain from it $\phi^{(t+1)}_{i,j}\forall(i,j)\in g_{1}$
end for
Return: ${\bm{\phi}^{(0)},\bm{\phi}^{(1)},...,\bm{\phi}^{(Num)}}$
Algorithm 1 MCMC sampling procedure using PBMG-XY
3 Application to the $\phi^{4}$ Theory
The lattice $\phi^{4}$ theory on a 2D lattice has undergone extensive investigation in both statistical mechanics and quantum field theory due to its intricate phase structure and critical behaviour. It demonstrates second-order phase transitions, characterized by alterations in system properties like magnetization or correlation length with varying coupling parameters. For a thorough comprehension of the lattice $\phi^{4}$ model, one can refer to the works [46, 47, 48]. In the following section, we will discuss the proposed ML-based sampling approach for the lattice $\phi^{4}$ theory.
3.1 Target Distribution
The Euclidean action for scalar $\phi^{4}$ theory in 2D can be written as,
$S(\bm{\phi},\lambda,m^{2})=\sum_{i,j}\Bigl{(}m^{2}+4\Bigr{)}\phi^{2}_{i,j}-\phi_{i,j}[\phi_{i+1,j}+\phi_{i,j+1}+\phi_{i-1,j}+\phi_{i,j-1}]+\lambda\phi_{i,j}^{4}$ (14)
where $(i,j)$ represents the indices of a lattice site and $\phi_{i,j}$ is a real-valued random variable defined for every lattice site.
The lattice distribution is given by the Boltzmann distribution law,
$p({\bm{\phi}|\lambda,m^{2}})\propto e^{-S(\bm{\phi},\lambda,m^{2})}$ (15)
The local action for lattice site $(i,j)$ [i.e.,the Lagrangian density] can be written as
$S_{loc}(\phi_{i,j},\lambda,m^{2},\{\phi_{l,m}:(l,m)\in n(i,j)\})$ $=S_{loc}(\phi_{i,j},\lambda,m^{2},\kappa_{i,j})$ (16)
$=\Bigl{(}m^{2}+4\Bigr{)}\phi_{i,j}^{2}+\lambda\phi^{4}_{i,j}-2\phi_{i,j}\kappa_{i,j}$ (17)
where $\kappa_{i,j}=\phi_{i+1,j}+\phi_{i,j+1}+\phi_{i-1,j}+\phi_{i,j-1}$.
We see that the local action for a lattice site depends on its four nearest neighbours i.e. the fields at each lattice site only interact with its nearby lattice points. Therefore, the conditional distribution of the lattice site $(i,j)$ can be written as
$p\big{(}\phi_{i,j}|\lambda,m^{2},\kappa_{i,j})=p\big{(}\phi_{i,j}|\mathbf{v}_{i,j})\propto e^{-S_{loc}(\phi_{i,j},\mathbf{v}_{i,j})}$ (18)
where, $\mathbf{v}_{i,j}=(\lambda,m^{2},\kappa_{i,j})$ is the condition vector for the distribution. Eq. 18 represents the target distribution which we model using the proposed method. For $\phi^{4}$ theory as well, we have divided the lattice into the same two partitions $g_{0}$ and $g_{1}$, where
$g_{k}=\{(i,j):(i+j)\%2=k\},k=0,1$ (19)
3.2 Modeling the Proposal Distribution: PBMG-$\phi^{4}$
We construct the proposal distribution for the scalar $\phi^{4}$ theory by using a Gaussian Mixture Model with six Gaussian components. The proposal distribution parameterized by $\bm{\theta}$, can be written as
$q(\phi_{i,j}|\mathbf{v}_{i,j};\bm{\theta})=\sum_{k=1}^{6}\pi_{k}(\mathbf{v}_{i,j};\bm{\theta}_{k})\mathcal{N}(\phi_{i,j}|\mu_{k}(\mathbf{v}_{i,j};\bm{\theta}_{k}),\sigma_{k}(\mathbf{v}_{i,j};\bm{\theta}_{k}))$ (20)
where $\mu_{k}$, $\sigma_{k}$ $\pi_{k}$ are the mean, standard deviation and mixing coefficients of the $k^{th}$ Gaussian distribution, and $\bm{\theta}=\{\bm{\theta}_{k}\}_{k=1}^{6}$. The parameters $\mu_{k}$, $\sigma_{k}$ $\pi_{k}$ are a function of the condition vector $\mathbf{v}_{i,j}=(\lambda,m^{2},\{\phi_{l,m}:(l,m)\in n(i,j)\})$ through a neural network parametrized by $\bm{\theta}_{k}$ which are learnt using a suitable loss function. The input to the $k^{th}$ neural network is the condition vector $\mathbf{v}_{i,j}$, and the outputs are the parameters $(\mu_{k},\log(\sigma_{k}),\pi_{k})$.
The architectures of all six neural networks are the same, with the only difference lying in the initialization of the network parameters. Each neural network consists of one hidden layer with 500 neurons and a ReLU activation function. The neurons in the final layer use linear activation. The value of $\log(\sigma_{k})>1$ is clipped to $1$. Since $\sum_{k}\pi_{k}=1$, the networks output logit values that are converted to $\pi_{k}$ by applying softmax.
3.3 Training and Inference Procedure
[…]
$\mathcal{L}=\mathbb{E}_{\mathbf{v}_{i,j}\sim p_{v}(\mathbf{v}_{i,j})}\left[\mathbb{E}_{\phi_{i,j}\sim q(\phi_{i,j}|\mathbf{v}_{i,j};\bm{\theta})}\left[\log{\frac{q(\phi_{i,j}|\mathbf{v}_{i,j};\bm{\theta})}{p\big{(}\phi_{i,j}|\mathbf{v}_{i,j})}}\right]+\|\bm{\pi}(\mathbf{v}_{i,j};\bm{\theta})\|\right]$ (21)
And the Monte Carlo approximation to the above expression is
$\mathcal{L}\approx\frac{1}{n}\sum_{r=1}^{n}\left[\frac{1}{N}\sum_{k=1}^{N}\left[\log q((\phi_{i,j})_{k}|(\mathbf{v}_{i,j})_{r};\bm{\theta})-\log p\big{(}(\phi_{i,j})_{k}|(\mathbf{v}_{i,j})_{r})\right]+\|\bm{\pi}\left((\mathbf{v}_{i,j})_{r};\bm{\theta}\right)\|\right]$ (22)
We sample from $p_{v}(\mathbf{v}_{i,j})$ to generate 17,500 samples of $\mathbf{v}_{i,j}$ and use this as the training set. We use the Adam optimizer with a learning rate of 0.0001 and default hyperparameters. We perform validation by calculating the average acceptance rate for 50 random sets of parameters ($\lambda$, $m^{2}$) where $\lambda\in[2.5,15]$ and $m^{2}\in[-8,0]$, and stop training when we achieve an average acceptance rate of around 98%. PBMG-$\phi^{4}$ also gets trained very quickly taking only a few minutes to train on a low-end single GPU machine.
The procedure for MCMC sampling using PBMG-$\phi^{4}$ is exactly the same as that of PBMG-XY given in Algorithm-2. During inference, however, while calculating $V_{g_{k}}$ where $k=0,1$, we take the absolute value of $\kappa_{i,j}$ and sample $\phi_{i,j}$ from the resulting GMM. We then consider $-\phi_{i,j}$ as the final sampled value and proceed with inference. This procedure is correct because,
$S_{loc}(\phi_{i,j},\lambda,m^{2},-\kappa_{i,j})=S_{loc}(-\phi_{i,j},\lambda,m^{2},\kappa_{i,j})$ (23)
4 Experiments and Results for PBMG-XY
[…]
Mean Energy: The Hamiltonian of a lattice sample divided by the number of lattice points is the mean energy $E(\bm{\phi}_{i})$ of that lattice sample $\bm{\phi}_{i}$. The observable mean energy is defined as the average mean energy per lattice sample.
$\langle E\rangle=\frac{1}{N}\sum_{i=1}^{N}E(\bm{\phi}_{i})=\frac{1}{N}\cdot\frac{1}{L^{2}}\sum_{i=1}^{N}H(\bm{\phi}_{i})$ (24)
Earth Mover Distance: EMD (aka Wasserstein Metric) is a distance metric and denotes, in a sense, the distance between any two probability distributions. The mathematical definition of EMD is given below. Here, $H_{G}$ and $H_{M}$ are the normalized histograms generated out of the ground truth samples and the samples from PBMG-XY respectively.
$\text{EMD}(H_{M},H_{G})=\sum_{x=-\infty}^{+\infty}\left|\sum_{t=-\infty}^{x}\left(H_{M}(t)-H_{G}(t)\right)\right|$ (25)
Percentage Overlap (%OL): %OL is a similarity metric and, as the name suggests, gives the overlap (in percentage) between the two normalized histograms $H_{M}$ and $H_{G}$ with bins indexed with $i$. The formula for %OL is as follows.
%$\text{\%OL}(H_{M},H_{G})=\sum_{i}\min{(H_{M}(i),H_{G}(i))}$ (26)
The plain-MCMC algorithm is used for the generation of ground truth. Different ensemble sizes are generated across the 23 temperatures for each lattice size. These ensemble sizes are determined based on the expected auto-correlation levels. A measure of the autocorrelation level in a Markov chain is the integrated autocorrelation time. The definition of integrated autocorrelation time that we use is the following:
$\tau_{int}=\frac{1}{2}+\sum_{k}\frac{C(k)}{C(0)},$ (27)
Here, $C(k)$ is the auto-correlation function of the chain at lag $k$ and $C(0)$ is the auto-correlation at zero lag. Higher auto-correlation necessitates more samples to ensure accurate estimation of observables. After generating the ensemble, we carry out thinning i.e., select every $k$th sample, where $k$ depends on the amount of auto-correlation in the chain, resulting in a chain with an effective sample size (ESS) of 2000. Here, ESS = $N/(2\tau_{int})$. The relative precision and confidence level in estimation for an ESS of 2000 are 91.235% and 95% respectively according to calculations by Vats et al [49]. Here, burn-in is unnecessary as we initiate the Markov chain from a high probability region with all spins at zero.
In our PBMG-XY simulations too, different ensemble sizes are generated across the 23 temperatures for each lattice size based on the expected auto-correlation levels, but the final chain used to estimate observables has an ESS of 2000. Appendix D provides the details of the ensemble size generated for each temperature and lattice size and the value of $k$ for each of them for both plain-MCMC and PBMG-XY.
Figure 3: Mean Energy Plots superimposed for a) L=8 and b) L = 16.
Figure 4: Mean Energy Plots superimposed for a) L=32 and b) L = 64.
The plots of mean energy vs. temperature for all the lattice sizes are given in Figures 3 and 4. We observe that the graphs of the observables are matching to a great extent. The acceptance rate for PBMG-XY varies considerably but we observe an average acceptance rate of 80% over the 23 temperatures we conduct our experiments on.
Metric Lattice size PBMG-XY ImplicitGAN-T
EMD $\downarrow$ 8x8 0.0042 $\pm$ 0.0036 0.022 $\pm$ 0.012
16x16 0.0027 $\pm$ 0.0026 0.010 $\pm$ 0.005
32x32 0.0027 $\pm$ 0.0034 N/A
64x64 0.0013 $\pm$ 0.0014 N/A
%OL $\uparrow$ 8x8 75.795 $\pm$ 11.424 68.28 $\pm$ 20.72
16x16 80.581 $\pm$ 8.784 65.83 $\pm$ 18.35
32x32 71.981 $\pm$ 12.556 N/A
64x64 77.773 $\pm$ 10.714 N/A
Table 1: Metrics for Mean Energy
Table 1 shows the EMD and %OL values for the mean energy observable. The values reported are the mean values along with standard deviations of EMD and %OL across the 23 temperatures. We observe that the EMD between the ground truth and the model is quite low and the %OL between the two is quite high. We compare the results with the work done by Singh et. al [34] for the two lattice sizes that they conducted their experiments on, i.e., $L=8$ and $L=16$. We observe that the metric values produced by our model are better than the ImplicitGAN-T model proposed by Singh et al. All of this supports the high accuracy of our approach.
5 Experiments and Results for PBMG-$\phi^{4}$
In this section, we evaluate the PBMG-$\phi^{4}$ model against the ground truth which we generate using Hamiltonian Monte Carlo simulation. For the $\phi^{4}$ theory, we conduct two kinds of experiments. One is on the computation of the lattice observables, and the other is on the computation of integrated auto-correlation time.
5.1 Experiment on Observables
We compute various observables on the lattice ensembles such as the mean $\phi^{2}$ value, two-point susceptibility, and two-point correlation function from both HMC and PBMG-$\phi^{4}$.
One of the key observables estimated in lattice $\phi^{4}$ theory is the “correlation function” or “two-point function”. This observable measures the correlation or interaction between field operators at different lattice sites across space. The correlation function can be written as
$G_{c}(i,j)=\frac{1}{V}\sum_{l,m}[\langle\phi_{l,m}\phi_{i+l,j+m}\rangle-\langle\phi_{l,m}\rangle\langle\phi_{i+l,j+m}\rangle].$
Here, $V$ is the lattice volume. The zero momentum correlation function is defined as
$C(t)=\sum_{i}G_{c}(i,t),$
where $t=j$ is the time axis. The pole mass and two-point susceptibility can be derived from $C(t)$ and $G(i,j)$ ,
$m_{p}=\log\left[\frac{C(t+1)}{C(t)}\right];~{}\chi=\sum_{i}G(i,j).$
The mean magnetization can be estimated as
$\langle|\tilde{\phi}|\rangle:\\ \tilde{\phi}=\frac{1}{V}\sum_{i,j}\phi_{i,j}.$
The experiment on the observables has been carried out for six lattice sizes, i.e., $L=8,16,24,32,48,\text{and}~{}64$. The parameter $m^{2}$ is fixed at $-4$ for all the ensembles. While preparing the ground truth, the numbers of molecular dynamics (MD) steps $N$ and step-size $\epsilon$ in HMC are adjusted so that the acceptance rate and the auto-correlation time remain almost constant for all the lattice sizes in the experiment. For every lattice size, 50K samples are generated. We consider every fifth lattice sample for the calculation of observables, i.e., 10K samples are considered for the final estimation of each observable.
Figure 5: a)Two-point susceptibility compared between HMC and PBMG-$\phi^{4}$ for L=$8$ and b)Absolute mean $\phi$ value for L=64
In the case of PBMG-$\phi^{4}$ model, for all the lattice sizes, i.e., $L=$ 8, 16, 32, 64 and 128, we generate 100K samples, and every 10th sample is used for measurement, bringing down the number of samples to 10K. In these experiments too, burn-in is unnecessary as we initiate the Markov chain from a high probability region with all field values equal to zero.
While estimating uncertainties, we use bootstrap error analysis with a bin size of $10$ for all the observables. In Figure 4(a), we calculate two-point susceptibility across the parameter $\lambda$ keeping $m^{2}=-4$ for $L=8$. Similarly, in Figure 4(b), we compute the absolute mean $\phi$ for larger lattice size, i.e. $L=64$. We can infer a second-order phase transition behaviour characteristic of lattice $\phi^{4}$ theory in 2D.
Figure 6: Two point Correlation function compared between HMC and PBMG-$\phi^{4}$ for lattice size a)8×$\times 8$ 8 and b) $32\times 32$
We also calculate the zero momentum correlation function for L = 8 and 32 for $\lambda=5.4$ and $m^{2}=-4$ displayed in Figure 5(a) and Figure 5(b) respectively. Here, we observe the expected behaviour i.e., we observe that the correlation between the sites decreases as we move to the centre of the lattice from the origin, and it is symmetric due to the periodicity in the lattice.
Ensemble S1 S2 S3 S4 S5
$m^{2}$4 -4 -4 -4 -4
L 16 24 32 48 64
$\lambda$ 8 6.3 5.6 5.0 4.8
$m_{p}L$ 12.80(2) 12.79(4) 12.82(5) 12.87(8) 12.81(3)
Table 2: Set of ensemble studied where the parameter $\lambda$ and L are varied while keeping $m_{p}L$ fixed at $\approx 12.8$.
Figure 7: Effective mass computed for parameter $\lambda=5.4$ and $L=32$ from lattice ensemble generated by HMC and PBMG-$\phi^{4}$ model.
We also compute pole mass from both HMC and PBMG-$\phi^{4}$ model by varying $\lambda$ and lattice size $L$ as shown in Table 2. We tune $\lambda$ and $L$ such that the pole mass remains constant, and at the same time, we move towards the critical region. The pole mass for ensemble $S3$ is shown in Figure 7 which we compute from both HMC and PBMG-$\phi^{4}$ generated samples. We ignore points in the middle of the time axis due to high noise which is a standard phenomenon in lattice field theory simulations as can be seen from the correlator plot in Figure 5(b) and is not specific to PBMG-$\phi^{4}$. We see a significant agreement between all the observables which we calculate from HMC and PBMG-$\phi^{4}$ model within statistical uncertainty.
We also calculate the Earth Mover Distance and %OL values between PBMG-$\phi^{4}$ and HMC for $\lambda=4.8$ and $m^{2}=-4$ for the observable absolute mean magnetization. The final chains used for calculations have an ESS of 2000. We use 40 bins for generating the histogram corresponding to each distribution. The results are given in Table 3. For PBMG-$\phi^{4}$ too, we observe that the EMD between the ground truth and the model is quite low and the %OL between the two is quite high. We also compare these metrics for $L=8$ with the RKLD-based NF approach proposed by Albergo et. al [50] and observe that PBMG-$\phi^{4}$ is either at par or better than their model.
Metric Lattice size PBMG-$\phi^{4}$ RKLD NF
EMD $\downarrow$ 8x8 0.0008 0.0008
16x16 0.0030 N/A
32x32 0.0085 N/A
64x64 0.0081 N/A
%OL $\uparrow$ 8x8 90.206 89.949
16x16 88.786 N/A
32x32 96.614 N/A
64x64 89.940 N/A
Table 3: Metrics for Absolute Mean Magnetization for PBMG-$\phi^{4}$
5.2 Experiment on Autocorrelation
Figure 8: Integrated autocorrelation time $\tau$ of two point susceptibility for lattice ensemble generated from HMC and PBMG-$\phi^{4}$
In HMC, the simulation cost rapidly increases as the volume of the lattice increases. Hence, simulation near the critical region becomes difficult, where one needs to generate larger ensembles due to the high correlation between samples, and also finer lattices with large lattice volumes. This correlation can be quantified by the integrated autocorrelation time.
To compare the efficiency of our model for the simulation of larger lattices, we compute integrated auto-correlation time ($\tau_{int}$) for different lattice sizes from $8\times 8$ upto $400\times 400$ from both HMC and PBMG-$\phi^{4}$ for the observable two-point susceptibility. We perform these experiments for $m^{2}=-4$ and $\lambda=5.4$, such that we remain quite close to the critical region of the theory. For each lattice volume, 10k samples are generated from both HMC and PBMG-$\phi^{4}$ model. For HMC, we fixed the parameter MD step size to $0.01$ and the number of MD steps to $20$ so that for the smallest lattice, the acceptance rate and $\tau_{int}$ are comparable for both methods.
Figure 8 shows how $\tau_{int}$ varies as we increase the lattice size in both methods. As we sample for larger lattice sizes, $\tau_{int}$ increases rapidly in the case of HMC. On the other hand, from the PBMG-$\phi^{4}$ model, we see a constant $\tau_{int}$ as we sample for different lattice sizes. Furthermore, we observe an almost-constant acceptance rate of around 98% from our model PBMG-$\phi^{4}$ irrespective of the lattice size and the parameters for which the simulation is carried out. On the other hand, from HMC, the sampling cost increases rapidly for generating larger lattices with an acceptance rate as high as that. This is a significant gain in terms of sampling cost while simulating a larger lattice volume. The detailed cost analysis is presented in the next section.
6 Computational Cost Analysis
In this section, we carry out the computational cost analysis of the PBMG approach for the two applications. We carry out this analysis for a particular set of values of the parameters ($T$ for the XY-model and $(\lambda,m^{2})$ for $\phi^{4}$ theory) and across different lattice sizes. We compare our algorithms with the baseline methods on the basis of the amount of time taken to generate roughly one independent sample $T_{eff}$. This quantity is equal to the amount of time taken to generate one sample ($T_{0}$) multiplied by two times the integrated autocorrelation time ($\tau_{int}$). Therefore,
$T_{eff}=T_{0}\cdot 2\tau_{int}.$ (28)
Here, $T_{0}$ is calculated for each lattice size by taking an average over 10 runs wherein each run 1000 samples are generated.
$\Delta(T_{eff})=\Delta(T_{0})\Delta(\tau_{int}).$ (29)
Here, $\Delta(x)$ denotes the ratio of quantity $x$ for the proposed model to the quantity $x$ for the baseline model. Smaller the value of $\Delta(T_{eff})$, the better the performance of the proposed model over the baseline model, in terms of the computational cost.
Many approaches [13, 39, 14, 21, 40, 41] either require re-training of proposed models for every new set of parameters, or train the proposed models for a particular lattice size and carry out inference on the same size. With our proposed approach, however, the same model, once trained, can be used for different parameters and lattice sizes. This leads to a significant reduction in computational costs. Below we discuss the computational costs during inference. The inference from the PBMG models is carried out on a low-end single GPU machine.
6.1 Analysis for PBMG-XY
The cost analysis for PBMG-XY is carried out for $T=1.0065$ which is very close to the critical temperature. We calculate $\tau_{int}$ for $T=1.0065$ by running chains of sufficient length for each lattice size similar to what is done in section 5.2. The baseline here is Plain-MCMC. The cost analysis for PBMG-XY is given in Table-4.
L 8 16 32 64 128 256 400
[…]
$\Delta(T_{0})$ 4.769 1.234 0.285 0.075 0.018 0.0048 0.0021
$\tau_{int}$ PBMG-XY 2.919 3.679 4.133 7.483 17.382 56.855 100.958
$\tau_{int}$ Plain-MCMC 3.980 9.393 18.972 65.175 89.282 84.67 86.792
$\Delta(\tau_{int})$ 0.733 0.392 0.218 0.115 0.195 0.671 1.163
$\Delta(T_{eff})$ 3.496 0.484 0.062 0.0086 0.0035 0.0032 0.0024
Table 4: Cost Analysis for PBMG-XY
We see that PBMG-XY is faster than Plain-MCMC from $L=16$ onwards, with the former being faster than the latter by 16 times for $L=32$ and by 416 times for $L=400$. $\Delta(T_{eff})$ consistently decreases, thereby becoming better, with an increase in lattice size. From these observations, we can conclude that PBMG-XY is faster than plain-MCMC for large lattice sizes.
6.2 Analysis for PBMG-$\phi^{4}$
The cost analysis for PBMG-$\phi^{4}$ is carried out for $m^{2}=-4$ and $\lambda=5.4$ which is very close to the critical region. The values $\tau_{int}$ produced in section 5.2 are used here for cost analysis. The baseline here is HMC. The cost analysis for PBMG-$\phi^{4}$ is given in Table-5.
L 8 16 32 64 128 256 400
[…]
$\Delta(T_{0})$ 5.143 3.474 2.677 2.833 1.098 0.274 0.065
$\tau_{int}$ PBMG-$\phi^{4}$ 6.136 9.918 8.547 8.803 9.1348 9.4632 8.881
$\tau_{int}$ HMC 1.85 2.547 3.417 13.457 20.0234 85.042 202.021
$\Delta(\tau_{int})$ 3.317 3.894 2.501 0.654 0.456 0.1112 0.0439
$\Delta(T_{eff})$ 17.059 13.528 6.695 1.853 0.501 0.030 0.0029
Table 5: Cost Analysis for PBMG-$\phi^{4}$
We see that PBMG-$\phi^{4}$ faster than HMC from $L=128$ onwards, with the former being faster than the latter by 33 times for $L=256$ and by 344 times for $L=400$. Here too, $\Delta(T_{eff})$ consistently decreases, thereby becoming better, with an increase in lattice size. From these observations, we can conclude that PBMG-$\phi^{4}$ is faster than HMC for large lattice sizes.
We conducted experiments till lattice size $L=8000$ for both PBMG-XY and PBMG-$\phi^{4}$ and witnessed a value of $T_{0}\approx 0.011s$ and $T_{0}\approx 0.012s$ respectively, which is almost the same as that for $L=8$ from the respective models. From these observations, we conclude that the time taken to generate one sample from the PBMG models is independent of the lattice size at least till $L=8000$ which is a considerable advantage over HMC and Plain-MCMC which, very clearly, take large amounts of time for large lattice sizes.
7 Conclusion
In this study, we introduced a novel approach, Parallelizable Block Metropolis-within-Gibbs (PBMG), designed to address the challenge of efficiently sampling Boltzmann distributions for large lattices. By decomposing the complicated joint distribution into simpler local conditional distributions, our approach leverages the power of generative machine learning, offering a scalable solution that transcends the limitations of traditional sampling methods. Through our experiments on the XY model and the Scalar $\phi^{4}$ theory, we have demonstrated the correctness and efficiency of our approach.
The PBMG models are trained using the Reverse KL objective, negating the need for existing simulated data, and get trained quickly taking less than an hour for PBMG-XY and only a few minutes for PBMG-$\phi^{4}$. The trained models can be employed across various lattice sizes and are conditioned on the action parameters enabling us to simulate using them for various lattice sizes and parameter values. The design of the proposal ensures that the acceptance rate remains unaffected by the lattice size under simulation.
For both PBMG-XY and PBMG-$\phi^{4}$, the results exhibit a significant match of observables and commendable values of metrics, providing a solid foundation for the accuracy of our approach. Both PBMG-XY and PBMG-$\phi^{4}$ prove to be faster than the baselines for larger lattice sizes. Furthermore, PBMG-$\phi^{4}$ showcases very high and consistent acceptance rates, which forms the basis for efficient sampling from it. The most remarkable observation from our experiments is the constant integrated autocorrelation time across different lattice sizes exhibited by PBMG-$\phi^{4}$, a clear testament to its efficiency in sampling and capturing physics near the critical region. We also analysed the amount of computational advantage that the PBMG models produce with both PBMG-XY and PBMG-$\phi^{4}$ exhibiting several times faster runtime for larger lattices.
PBMG is easily extensible to 3D and 4D lattices, which once achieved, would considerably ease simulating large 3D and 4D lattices for analysis. Additionally, extending PBMG for non-local models is another direction of research which holds promising returns.
    A.1 Probabilistic formulation of the XY model
    The XY model is one of the simpler lattice models in statistical mechanics. Each lattice site in this model consists of a vector of unit magnitude $\bm{s}_{i,j}=(\cos\phi_{i,j},\sin\phi_{i,j})^{T}$ where $\phi_{i,j}\in[0,2\pi)$ and $(i,j)$ denote the exact position of the lattice site. These unit vectors are referred to as spins. A spin $\bm{s}_{i,j}$ is characterised by the angle $\bm{\phi}_{i,j}$. The Hamiltonian of the configuration $\bm{\phi}$ is given by,
    $H(\bm{\phi})=-\sum_{\langle(i,j),(k,l)\rangle}\bm{s}_{i,j}\cdot\bm{s}_{k,l}=-\sum_{\langle(i,j),(k,l)\rangle}\cos(\phi_{i,j}-\phi_{k,l})$ (30)
    where the sum is taken such that the lattice sites $(i,j)$ and $(k,l)$ are neighbors. Each pair of neighbours is counted exactly once. Here, we have considered only those nearest-neighbour interactions which are ferromagnetic in nature i.e., we have assumed $J>0$. Furthermore, we have taken $J\equiv 1$ i.e., we’ve taken it as the unit of energy.
    Using the Boltzmann distribution formulation, the probability density function for a temperature $T>0$ is,
    $p(\bm{\phi};T)=\frac{1}{Z(T)}\exp\left({-\frac{H(\bm{\phi})}{T}}\right)$ (31)
    Here, the normalizing constant $Z(T)$ (also known as partition function) is the integral of $\exp\left(-\frac{H(\bm{\phi})}{T}\right)$ over all possible configurations. Note that we have set the Boltzmann constant($k_{B}$) to unity over here.
    A.2 Observables in the 2D XY model
    The macroscopic properties of the XY model at a given temperature that can be measured or observed are referred to as observables. From a probabilistic perspective, these observables are actually expected values of certain lattice functions(by this we mean functions that take in a lattice as input and give some output) over all possible configurations of lattices. Mathematically,
[…]
    A.3 The BKT transition
    ‘BKT’ in the BKT transition stands for Berezinskii-Kosterlitz-Thouless. This transition is a characteristic phenomenon of the XY model. It is a phase transition that takes place at the critical temperature $T_{c}$ due to the unbinding of vortex-antivortex pairs. It is a first-order transition and the magnetic susceptibility(and other parameters of similar order) diverge because of it. This transition is the cause behind the relatively difficult simulation of the XY model as compared to other models with a higher-order parameter that smoothly changes with temperature. Simulation of the XY model in the critical region(i.e. near the critical temperature $T_{c}$) is relatively difficult as compared to simulation at lower or higher temperatures.
    Appendix B Baseline algorithm for experiments on the XY model
    The procedure for MCMC sampling using the plain-MCMC algorithm for the XY model is briefed below.
    Initialize: $\bm{\phi}$ as $\bm{\phi}^{(0)}$
    for $t=0$ to $Num-1$ do
           for $m=0$ to $L^{2}$ do
[…]
                 Propose a new spin value at $(x,y)$ by sampling from Unif[0, $2\pi$).
                 Calculate the change in energy.
                 Calculate the acceptance probability using the change in energy. Update the lattice accordingly.
           end for
          Save the lattice as $\bm{\phi}^{(t)}$.
    end for
    Return: ${\bm{\phi}^{(0)},\bm{\phi}^{(1)},...,\bm{\phi}^{(Num)}}$
    Algorithm 2 Plain-MCMC sampling algorithm in our experiments
    The plain-MCMC algorithm consists of only a single M-H accept-reject step between successive lattices. The only difference in the way we have implemented it is that, we have taken a default thinning value of $k=L^{2}$ i.e., successive lattices differ by $L^{2}$ M-H accept-reject steps. In this way, we would have subsequent lattices that are considerably different from each other whereas in the original algorithm, only one lattice site would be different and that too if the proposed value gets accepted.
    Our implementation makes the autocorrelation time existing in the Markov chain seem more reasonable. For example, as reported in Appendix D, for a 64x64 lattice, the value of $2\tau_{int}$ for $T=1.0065$ is 131. If the original implementation had been used, this value of $2\tau_{int}$ would have been around $64*64*131=536,576$. In comparison to 536576, 131 is more reasonable to work with and is of the same order as that of $2\tau_{int}$ for PBMG-XY which is 10. Note that, in the original implementation, $2\tau_{int}$ would be almost $L^{2}$ times higher, but $T_{0}$ would also be $L^{2}$ times smaller, keeping the values of $T_{eff}$ and, therefore, $\Delta(T_{eff})$ very nearly the same.
    Appendix C Metrics for experiments on the XY model
    We discuss here the different metrics that we have used to quantify the performance of our model. These metrics, essentially, indicate the similarity between two histograms or probability distributions and we use these metrics because we want to estimate the similarity between the histogram we generate out of our model and the histogram generated from plain-MCMC.
    C.1 Earth Mover Distance
    Considering each distribution to be a pile of dirt, EMD is the minimum cost of transforming one pile into the other. A lower value of EMD indicates a higher matching percentage of the two probability distributions. For each lattice size(L) and for each temperature, we generate a number of lattices from the model and from plain-MCMC(which is taken as the ground truth). We then generate normalized histograms $(H_{M}\&H_{G})$ out of these two sets of lattices and then calculate the EMD between these two histograms.
    $\text{EMD}(H_{M},H_{G})=\sum_{x=-\infty}^{+\infty}\left|\sum_{t=-\infty}^{x}\left(H_{M}(t)-H_{G}(t)\right)\right|$ (33)
    C.2 Percentage Overlap
    The accuracy of %OL is less than that of EMD because relatively accurate observables may look dissimilar w.r.t this metric i.e., may get a small value of percentage overlap. To calculate %OL we take the minimum of the two histogram values for each index and sum over all indices. In the expression below, $i$ is the bin index corresponding to which the minimum is being calculated.
    %$\text{\%OL}(H_{M},H_{G})=\sum_{i}\min{(H_{M}(i),H_{G}(i))}$ (34)
    Appendix D Details of experiments on the XY model
    Temperature 8x8 16x16 32x32 64x64
    S k S k S k S k
    0.137 36,000 18 66,000 33 74,000 37 66,000 33
    0.224 34,000 17 44,000 22 46,000 23 68,000 34
    0.311 34,000 17 48,000 24 42,000 21 54,000 27
    0.398 26,000 13 32,000 16 42,000 21 40,000 20
    0.485 18,000 9 28,000 14 40,000 20 40,000 20
    0.572 20,000 10 28,000 14 30,000 15 30,000 15
    0.659 26,000 13 22,000 11 28,000 14 34,000 17
    0.746 24,000 12 28,000 14 44,000 22 30,000 15
    0.833 20,000 10 42,000 21 40,000 20 46,000 23
    0.920 22,000 11 60,000 30 44,000 22 48,000 24
    1.0065 16,000 8 38,000 19 76,000 38 262,000 131
    1.093 48,000 24 142,000 71 206,000 103 138,000 69
    1.180 34,000 17 58,000 29 50,000 25 68,000 34
    1.267 46,000 23 48,000 24 50,000 25 48,000 24
    1.354 42,000 21 40,000 20 30,000 15 38,000 19
    1.441 26,000 13 30,000 15 24,000 12 26,000 13
    1.528 16,000 8 20,000 10 18,000 9 26,000 13
    1.615 18,000 9 18,000 9 18,000 9 20,000 10
    1.702 14,000 7 14,000 7 18,000 9 18,000 9
    1.789 14,000 7 12,000 6 16,000 8 18,000 9
    1.876 10,000 5 16,000 8 12,000 6 16,000 8
    1.963 8,000 4 10,000 5 12,000 6 12,000 6
    2.05 8,000 4 12,000 6 10,000 5 12,000 6
    Table 6: Ensembles sizes and value of $k$ for all temperatures and lattice sizes for ground truth generation from plain-MCMC
    Temperature 8x8 16x16 32x32 64x64
    S k S k S k S k
    0.137 18,000 9 20,000 10 14,000 7 22,000 11
    0.224 8,000 4 16,000 8 16,000 8 18,000 9
    0.311 22,000 11 14,000 7 16,000 8 26,000 13
    0.398 10,000 5 10,000 5 18,000 9 24,000 12
    0.485 6,000 3 10,000 5 26,000 13 24,000 12
    0.572 6,000 3 8,000 4 22,000 11 28,000 14
    0.659 6,000 3 16,000 8 20,000 10 18,000 9
    0.746 8,000 4 10,000 5 24,000 12 18,000 9
    0.833 10,000 5 8,000 4 20,000 10 22,000 11
    0.920 8,000 4 16,000 8 20,000 10 22,000 11
    1.0065 12,000 6 20,000 10 16,000 8 20,000 10
    1.093 16,000 8 64,000 32 18,000 9 16,000 8
    1.180 18,000 9 24,000 12 20,000 10 24,000 12
    1.267 28,000 14 16,000 8 24,000 12 24,000 12
    1.354 16,000 8 12,000 6 20,000 10 20,000 10
    1.441 8,000 4 10,000 5 16,000 8 24,000 12
    1.528 8,000 4 8,000 4 14,000 7 24,000 12
    1.615 6,000 3 6,000 3 22,000 11 22,000 11
    1.702 6,000 3 6,000 3 18,000 9 20,000 10
    1.789 6,000 3 6,000 3 20,000 10 24,000 12
    1.876 4,000 2 4,000 2 20,000 10 24,000 12
    1.963 6,000 3 4,000 2 16,000 8 20,000 10
    2.05 4,000 2 4,000 2 20,000 10 22,000 11
    Table 7: Ensembles sizes and value of $k$ for all temperatures and lattice sizes for data generated from PBMG-XY